# SPHEREx Pre-Perihelion Mapping of $\mathrm{H_{2}O}$, $\mathrm{CO_{2}}$, and $\mathrm{CO}$ in Interstellar Object 3I/ATLAS

### v2

@@ V Discussion
All of the observations reported here (the giant $\mathrm{CO}_{2}$ coma, the lack of any observed lightcurve variability, the prevalence of large coma grains, and the 100x greater coma scattered light flux than a $2.5\,\mathrm{km}$ radius, 0.04 geometric albedo spherical object would produce) are consistent with an object spectrally dominated by bright, $\mathrm{CO}_{2}$- and water ice-rich coma material. The low amount of $\mathrm{H_{2}O}$ gas production observed by SPHEREx in August 2025 can be explained by the object’s having only recently arrived from the cold depths of interstellar space and its remoteness from the Sun (outside the water ice line between $3.4$ and $3.1\,\mathrm{au}$) during the observational period (Lisse et al., 2021, 2022a). The lack of strong $\mathrm{CO}$ emission is slightly surprising, as the $\mathrm{CO}$-ice line in the solar system lies at $\sim 40\,\mathrm{au}$ from the Sun, until one realizes it is an indication of hypervolatile depletion.
@@
A strong predominance of $\mathrm{CO}_{2}$ gas emission coupled with $\mathrm{CO}$ hypervolatile depletion has been seen in solar system hyperactive comets like 103P (A’Hearn et al., 2012, $Q_{\mathrm{CO}_{2}}/Q_{\mathrm{H_{2}O}}\sim 20\,\mathrm{\%}$; $Q_{\mathrm{CO}}/Q_{\mathrm{H_{2}O}}\sim 0.15\,\mathrm{\%}$) and 46P/Wirtanen (McKay et al., 2021, $Q_{\mathrm{CO}_{2}}/Q_{\mathrm{H_{2}O}}\sim 15\,\mathrm{\%}$; $Q_{\mathrm{CO}}/Q_{\mathrm{H_{2}O}}<0.5\,\mathrm{\%}$). In fact, 103P is the comet with the lowest known ratio of $\mathrm{CO}$ to $\mathrm{CO}_{2}$ (Harrington Pinto et al., 2022, see Fig. 9a). Hyperactive comets are defined as comets with gas production rates many times what sublimation from their nucleus surfaces can/should be able to support; their gas production is instead produced primarily from large, long-lived icy grains emitted into the surrounding coma by the nucleus.
@@
Figure 4: (a) $\mathrm{CO}/\mathrm{CO}_{2}$ gas production rate ratios of solar system comets (Harrington Pinto et al., 2022). The $\mathrm{CO}$ or $\mathrm{CO}_{2}$ dominance criteria of $1\pm 0.33$ is indicated as horizontal shadow (Harrington Pinto et al., 2022). The ratio of 3I makes it extremely $\mathrm{CO}$ poor and equivalent to the most thermally processed comet in the 2022 survey, 103P. (b) Pre-perihelion optical lightcurves observed for comet 103P during the 2010–2011 apparition and 3Iin May - Oct 2025. Published data from (Jewitt & Luu, 2025; Tonry et al., 2025; Ye et al., 2025) have been shifted to match the absolute magnitude of Minor Planet Center (MPC) Database [^7] https://www.minorplanetcenter.net/db_search/show_object?object_id=3I for 3I and https://www.minorplanetcenter.net/db_search/show_object?object_id=103P for 103P. , the T-mag from JPL Horizons ephemerides service [^8] https://ssd.jpl.nasa.gov/horizons/app.html, and the data kindly shared by S. Yoshida (2025999http://www.aerith.net/comet/catalog/0003I/2025N1.html, priv. comm.). There is very good agreement between the two inbound lightcurve trends. An inherent spread of $\sim 0.5\,\mathrm{mag}$ in the measurements between nights and between observers can be seen. (c) Close flyby imaging of the $\sim 1.5\times 0.6\times 0.6\,\mathrm{km}$ comet 103P nucleus as obtained by the Deep Impact Extended mission (A’Hearn et al., 2011). A marked bilobate nuclear morphology was seen, with unusually large ($m$ to dm) chunks of $\mathrm{CO}_{2}$ rich ice and gas emanating from the end of the smaller lobe and the middle of the larger lobe. The surrounding coma, filled with large long lived icy grains, was found to be at least 5 times brighter and more actively emitting gas than the nucleus. While the nucleus of 3I has never been resolved, given its similar behavior, it could possibly share similar morphological characteristics. (d) Schematic of the 3I object derived from SPHEREx $\mathrm{CO}_{2}$ photometric mapping. The $2.5\,\mathrm{km}$ radius nucleus (Jewitt et al., 2025) was not resolved at the SPHEREx pixel scale of $6.15\arcsec\approx 11,600\,\mathrm{km}$ at $2.6\,\mathrm{au}$ distance, but a notional nucleus image has been included at the origin to give the reader its location and a sense of scale. Two distances have been marked off: The $\sim 38,000\,\mathrm{km}$ (white dashed circle) inside of which we think coma ice is actively sourcing $\mathrm{CO}_{2}$ by sublimation; and $\sim 340,000\,\mathrm{km}$ (orange dashed oval), the minimum radial distance to which we detect $\mathrm{CO}_{2}$ gas at 2-$\sigma$ level.
@@
103P was especially well studied during the close ($\sim 1000\,\mathrm{km}$ minimum distance) flyby of the NASA Deep Impact Extended mission (A’Hearn et al., 2011, in November 2010), which imaged large chunks of $\mathrm{CO}_{2}$ rich material being ejected “in a reverse snow” from one of the comet’s lobes and measured a coma $\sim 5\times$ brighter than the nucleus in total scattered light flux (Figure 4c). Because of 103P’s small size but near typical gas production rate, it has been argued to be near the end of its life as a comet in the inner solar system (Lisse et al., 2009). The large chunks of $\mathrm{CO}_{2}$-rich material are explained as primordial remnants of solid $\mathrm{CO}_{2}$ ice phases protected down deep near the center of an originally much larger nuclear body which has lost its much more hypervolatile ice phases (e.g., $\mathrm{CO}$, $\mathrm{N_{2}}$, and $\mathrm{CH_{4}}$). The imaging results from the Deep Impact extended mission suggest that the regions of its nucleus emitting $\mathrm{CO}_{2}$-rich ice are separated from regions producing water gas, and the source ices may be physically separated.
V.1 Other Hypotheses
@@
There are several other hypotheses that have been proposed to explain 3I’s physical behavior. For example, it has been proposed that 3I originated in a very $\mathrm{CO}$-poor but $\mathrm{CO}_{2}$-rich exosystem (i.e., a/the nature hypothesis to our evolutionary nurture proposal of advanced thermal evolution). We find this explanation problematic, as this would imply a system with a host star C/O ratio <1/10<1/10 the Sun’s, and a dearth of reduced aliphatic and aromatic species as well. [C/Fe] and [C/H] <1.0<-1.0 for a nearby Milky Way star is highly improbable (Hinkel et al., 2017; Lisse et al., 2017; Mészáros et al., 2025)], and the aliphatic and PAH absorption features for 3I in the JWST NIRSPEC spectrum of Cordiner et al. (2025) look to be of reasonably normal strength.
It has also been proposed that 3I was sourced from an exosystem with near-solar composition, and it began as the same; but that galactic cosmic-rays preferentially transformed its near-surface $\mathrm{CO}$ ices into $\mathrm{CO}_{2}$ species via radiolytic driven production of OH, H, and O and subsequent oxidation of all $\mathrm{CO}$ to $\mathrm{CO}_{2}$. This proposed process would also tend to make $\mathrm{H_{2}CO}$, $\mathrm{CH_{3}OH}$, $\mathrm{HCOOH}$, as well, which may be consistent with current observations of 3I. However it is not consistent with the work of Moore et al. (2003), which showed that particle irradiation of either pure $\mathrm{CO}$ or $\mathrm{CO}_{2}$ in water ice tends to produce a mixture of the two species with abundance ratio on the order of unity, nor is it consistent with the $\mathrm{CO}/\mathrm{CO}_{2}$ ratios found in solar system Oort cloud comets, which spend $99\,\mathrm{\%}$ of their lifetimes in the ISM galactic cosmic ray radiation field, and emit abundant $\mathrm{CO}$. Close passages of Oort cloud comets by a star (the Sun), which 3I may never have experienced once its surface cr alteration began, would serve to drive off $\mathrm{CO}$ preferentially vs $\mathrm{CO}_{2}$, yet the $\mathrm{CO}/\mathrm{CO}_{2}$ abundance ratio remains high in these objects (Harrington Pinto et al., 2022).
We know that 3I is undergoing thermally driven evolution, as evidenced by the rate at which it is emitting dust and gas is trending like $r_{\mathrm{hel}}^{-3.8}$ after correcting for any observer-3I geometry (Jewitt & Luu, 2025), faster than the $r_{\mathrm{hel}}^{-2}$ trend for energy received from the sun is changing. We also know it is thermally evolving as we have watched its dust coma plus tail morphology go from large particle, sunward flying grains at discovery outside the water ice line, to the current small particle dominated, antisolar dust grains, as 3I’s the water production rate has picked up markedly inside the water ice line. We therefore consider a thermal evolutionary hypothesis for its behavior as not only highly likely, but necessary to consider.
V.2 Home System Implications
@@
A heavily thermally processed state for 3I has a number of important implications. Firstly, by analogy to 103P’s behavior, it tells us that 3I has undergone 100s to 1,000s of close passages by its natal host/birth star (we reject the possibility that 3I has undergone 100’s to 1000’s of close passages of random galactic stars on its passsge to the solar system; this rate of close passage for a single object, even one as old as the galaxy, would imply a huge rate of interstellar object passages & detections in our solar system, which is not seen). These close thermally altering passages could have been accomplished either as a primordial body formed in-place near its birth system’s $\mathrm{CO}_{2}$ ice line (to ensure minimal CO but extensive labile $\mathrm{CO}_{2}$ ice incorporation) and quickly ejected soon after PPD disk clearing, or more classically, as in our own mature system, as an ice-rich KBO that became scattered into its inner system for a few hundred near-star orbits before ejection.
It also implies that 3I had a close encounter with a massive body that ejected it from its home system. Very simply, the maximum $\Delta v$ (velocity change) a close planetary flyby can impart is approximately the escape velocity for the planet. In our solar system, with $V_{\mathrm{esc}}\sim 25\,\mathrm{km/s}$ at $2.5\,\mathrm{au}$ from the Sun, only Jupiter ($V_{\mathrm{esc}}\sim 65\,\mathrm{km/sec}$), Saturn ($35.5\,\mathrm{km/sec}$), and barely Neptune ($25.5\,\mathrm{km/sec}$) could eject objects directly. Allowing for the possibility of a smaller mass birth star than the Sun definitely means that an inner planet as small as Uranus or Neptune could serve as an ejector, but this is about the minimum mass required because the water ice line retreats rapidly inward as the host star mass decreases ($V_{\mathrm{esc}}$ from the ice line to the ISM is actually larger for an M8V dwarfs like TRAPPIST-1, $\sim 34\,\mathrm{km/sec}$).
Objects commonly in close orbit around the host stars that are more massive than planets, like brown dwarfs, main sequence stars, AGB stars, white dwarfs, neutron stars, and black holes have a large enough $V_{\mathrm{esc}}$ to cause ejection from the ice line of a system, so rather than an inner massive planet, 3I could have come from a close-in binary system (Jackson et al. 2018, 2018MNRAS.478L..49J).
V.3 ISO Evolution
@@
Finally we end with one last piece of big picture scholarly speculation. Naive models of ISO creation would suggest the maximal number of ISOs created by a system should be when it hosts the most free flying small planetesimals that are dynamically hot, that is, when the system is very young, within the first few Myrs of the host stars formation. But this is also the era when a system’s protoplanetary disk is optically thick near its midplane and planetesimals in the densest part of the disk remain cold and thermally unprocessed. PPDs are seen to clear within $5\mathrm{-}10\,\mathrm{Myrs}$ of host star formation though (Rieke et al., 2005; Ben et al., 2025, and references therein), and the first heavy thermal processing of KBOs estimated to occur in $1\mathrm{-}40\,\mathrm{Myrs}$ (Davidsson, 2021; Steckloff et al., 2021; Lisse et al., 2021, 2022b; Parhi & Prialnik, 2023). Planetary migration and Oort Cloud formation involving small body ejection processes can take many 100’s of Myrs to complete (Ref from Dearth papers). We can thus argue plausibly that PPDs can clear and small body thermal evolutionary processes can occur more rapidly than dynamical processes can cool down and mature in a system, and we should see a mix of volatile rich, hypervolatile depleted, and totally devolatilized ejected objects as our ISOs. This is precisely what we have seen so far in highly-devolatilized 1I/Oumuamua, hypervolatile-rich 2I/Borisov, and well-thermally processed 3I.
@@ VI Conclusions and Future Work
[…]
@@
Registration and co-adding of dozens of scattered light continuum images at 1.0 - 1.5 um produced a high S/N ratio image consistent with a stellar point source of no extension. The same images were used to create a lightcurve, for which variability $\lesssim 15\,\mathrm{\%}$ was seen from beginning to end of the observing period. The absolute brightness of 3I at 1.0 - 1.5 um is consistent with a nucleus of effective radius Rnuc $\lesssim$ 2.5 km, as reported by HST (Jewitt+ 2025) surrounded by a 100 times brighter coma. The SPHEREx spectral continuum measured over 2 weeks is highly consistent with the continuum reported by JWST during its single visit to the object on 06 Aug 2025 (Cordiner et al. 2025). The 1.5-4.0 um continuum structure shows a strong feature commensurate with water ice + refractory organics absorption. The overall shape of this reflectance continuum is very similar to that reported for the cliff-type KBOs by Pinilla-Alonso, N., Brunetto, R., De Prá, M.N. et al. 2025, “A JWST/DiSCo-TNOs portrait of the primordial Solar System through its trans-Neptunian objects”, Nat Astron 9, 230.
@@
The August SPHERExobservations provided the first detection of strong $\mathrm{CO}_{2}$ emission from the comet into a giant extended surrounding coma, proving that $\mathrm{CO}_{2}$ sublimation from large icy coma grains is the main driver for the observed outgassing activity and explaining its extended morphology as the product of $\mathrm{CO}_{2}$ gas sublimation that was undetectable in the first reported UVIS characterizations of the object. The moderate observed water gas production rate can be explained as an object just coming in from the cold depths of interstellar space that was observed at rh = 3.2 au from the Sun outside the water ice line.
@@
In many ways the observed activity behavior of 3I, including its preponderance of $\mathrm{CO}_{2}$ emission, lack of $\mathrm{CO}$ output, small size, and predominance of large chunks of water-ice rich icy material in a flux-dominant coma is reminiscent of the behavior of short period comet 103P, target of the NASA Deep Impact extended mission in 2010 and a “hyperactive comet” near the end of its outgassing lifetime (A’Hearn et al., 2011; Lisse et al., 2009). Such an object would have undergone hundreds to thousands of close perihelion passages by its natal birth star inside the system’s water ice line, as 103P has by our Sun. Whether this was due to 3I forming within a few au of its birth star or being scattered into this region after formation but before ejection is not clear. In terms of ice incorporation and thermal processing mechanisms operating on solar system comets (Lisse et al., 2021, 2022a), 3I is closer in evolution to barely- or non-active 1I/Oumuamua than primitive, ice rich 2I/Borisov.
@@
Compared to the naïve expectation that solar systems shed the majority of their comets in their first few Myrs of existence when their small icy planetesimals are most abundant, most hypervolatile ice rich, and most dynamically hot, so that all ISOs should be like 2I/Borisov, this suggests that ISOs are thermally processed quicker than they dynamically relax. This hypothesis needs to be tested by population studies of many more ISOs than the 3 currently known, and hopefully many more will be studied by the SPHERExmission’s all-sky survey. Another inference that can be made about 3I’s home system is that it contains a companion body of mass MNeptune or larger within a few au of the host star; such a massive body is required to eject a body like 3I from a stellar system with host star of class M8V or larger.
@@
This paper relies on data obtained by the SPHEREx Observatory (operated by Caltech and JPL on behalf of the National Aeronautics Space Administration under contract 80GSFC18C0011) that is hosted by IPAC as part of the IRSA archive (10.26131/IRSA652 SPHEREx Quick Release Spectral Images - QR2). We are indebted to the SPHEREx project for obtaining additional pointings at 3I, approximately 100 percent more, than their nominal survey would have obtained less than 3 months after the start of the survey by testing and implementing a new commanding protocol. Part of the research described in this paper was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). The authors also acknowledge the Texas Advanced Computing Center (TACC)  [^10] http://www.tacc.utexas.edu at The University of Texas at Austin for providing computational resources that have contributed to the research results reported within this paper.

### v3

@@ V Discussion
All of the observations reported here (the giant $\mathrm{CO}_{2}$ coma, the lack of any observed lightcurve variability, the prevalence of large coma grains, and the 100 times greater coma scattered light flux than a $2.5\,\mathrm{km}$ radius, 0.04 geometric albedo spherical object would produce) are consistent with an object spectrally dominated by bright, $\mathrm{CO}_{2}$- and water ice-rich coma material. The low amount of $\mathrm{H_{2}O}$ gas production observed by SPHEREx in August 2025 can be explained by the object’s having only recently arrived from the cold depths of interstellar space and its remoteness from the Sun (outside the water ice line between $3.4$ and $3.1\,\mathrm{au}$) during the observational period (Lisse et al., 2021, 2022a). The lack of strong $\mathrm{CO}$ emission is slightly surprising, as the $\mathrm{CO}$-ice line in the solar system lies at $\sim 40\,\mathrm{au}$ from the Sun, until one realizes it is an indication of hypervolatile depletion.
@@
Given the current turbulent mix of information currently posted on 3I  it is important to state that there are known, studied natural solar system comets producing analogous behavior. A strong predominance of $\mathrm{CO}_{2}$ gas emission coupled with $\mathrm{CO}$ hypervolatile depletion has been seen in solar system hyperactive comets like 103P (A’Hearn et al., 2012, $Q_{\mathrm{CO}_{2}}/Q_{\mathrm{H_{2}O}}\sim 20\,\mathrm{\%}$; $Q_{\mathrm{CO}}/Q_{\mathrm{H_{2}O}}\sim 0.15\,\mathrm{\%}$) and 46P/Wirtanen (McKay et al., 2021, $Q_{\mathrm{CO}_{2}}/Q_{\mathrm{H_{2}O}}\sim 15\,\mathrm{\%}$; $Q_{\mathrm{CO}}/Q_{\mathrm{H_{2}O}}<0.5\,\mathrm{\%}$). In fact, 103P is the comet with the lowest known ratio of $\mathrm{CO}$ to $\mathrm{CO}_{2}$ (Harrington Pinto et al., 2022, see Fig. 9a). Hyperactive comets are defined as comets with gas production rates many times what sublimation from their nucleus surfaces can/should be able to support; their gas production is instead produced primarily from large, long-lived icy grains emitted into the surrounding coma by the nucleus.
@@
Figure 4: (a) $\mathrm{CO}/\mathrm{CO}_{2}$ gas production rate ratios of solar system comets (Harrington Pinto et al., 2022). The $\mathrm{CO}$ or $\mathrm{CO}_{2}$ dominance criteria of $1\pm 0.33$ is indicated as horizontal shadow (Harrington Pinto et al., 2022). The ratio of 3I makes it extremely $\mathrm{CO}$ poor and equivalent to the most thermally processed comet in the 2022 survey, 103P. (b) Pre-perihelion optical lightcurves observed for comet 103P during the 2010–2011 apparition and 3Iin May - Oct 2025. Published data from (Jewitt and Luu, 2025; Tonry et al., 2025; Ye et al., 2025) have been shifted to match the absolute magnitude of Minor Planet Center (MPC) Database [^7] https://www.minorplanetcenter.net/db_search/show_object?object_id=3I for 3I and https://www.minorplanetcenter.net/db_search/show_object?object_id=103P for 103P. , the T-mag from JPL Horizons ephemerides service [^8] https://ssd.jpl.nasa.gov/horizons/app.html, and the data kindly shared by S. Yoshida (2025999http://www.aerith.net/comet/catalog/0003I/2025N1.html, priv. comm.). There is very good agreement between the two inbound lightcurve trends. An inherent spread of $\sim 0.5\,\mathrm{mag}$ in the measurements between nights and between observers can be seen. (c) Close flyby imaging of the $\sim 1.5\times 0.6\times 0.6\,\mathrm{km}$ comet 103P nucleus as obtained by the Deep Impact Extended mission (A’Hearn et al., 2011). A marked bilobate nuclear morphology was seen, with unusually large ($m$ to dm) chunks of $\mathrm{CO}_{2}$ rich ice and gas emanating from the end of the smaller lobe and the middle of the larger lobe. The surrounding coma, filled with large long lived icy grains, was found to be at least 5 times brighter and more actively emitting gas than the nucleus. While the nucleus of 3I has never been resolved, given its similar behavior, it could possibly share similar morphological characteristics. (d) Schematic of the 3I object derived from SPHEREx $\mathrm{CO}_{2}$ photometric mapping. The $2.5\,\mathrm{km}$ radius nucleus (Jewitt et al., 2025) was not resolved at the SPHEREx pixel scale of $6.15\arcsec\approx 11,600\,\mathrm{km}$ at $2.6\,\mathrm{au}$ distance, but a notional nucleus image has been included at the origin to give the reader its location and a sense of scale. Two distances have been marked off: The $\sim 38,000\,\mathrm{km}$ (white dashed circle) inside of which we think coma ice is actively sourcing $\mathrm{CO}_{2}$ by sublimation; and $\sim 340,000\,\mathrm{km}$ (orange dashed oval), the minimum radial distance to which we detect $\mathrm{CO}_{2}$ gas at 2-$\sigma$ level.
@@
103P was especially well studied during the close ($\sim 1000\,\mathrm{km}$ minimum distance) flyby of the NASA Deep Impact Extended mission (A’Hearn et al., 2011, in November 2010), which imaged large chunks of $\mathrm{CO}_{2}$ rich material being ejected “in a reverse snow” from one of the comet’s lobes and measured a coma $\sim 5\times$ brighter than the nucleus in total scattered light flux (Figure 4c). The large chunks of $\mathrm{CO}_{2}$-rich material are explained as primordial remnants of solid $\mathrm{CO}_{2}$ ice which have lost much of their more hypervolatile ice phases (e.g., $\mathrm{CO}$, $\mathrm{N_{2}}$, and $\mathrm{CH_{4}}$). The imaging results from the Deep Impact extended mission suggest that the regions of its nucleus emitting $\mathrm{CO}_{2}$-rich ice are separated from regions producing water gas, and the source ices may be physically separated.
V.1 Physical Behavior Hypotheses
@@
There are several hypotheses that have been proposed to explain 3I’s physical behavior. For example, it has been proposed that 3I originated in a very $\mathrm{CO}$-poor but $\mathrm{CO}_{2}$-rich exosystem (i.e., the “nature hypothesis” to the following “evolutionary nurture proposals”). Assuming that the $\mathrm{CO}:\mathrm{CO}_{2}$ ratio is directly reflective of the natal system’s C:O abundance ratio (c.f., Bodewits et al., 2020), we find this explanation problematic, as this would imply a system with a host star C/O ratio <1/10<1/10 the Sun’s, and a dearth of reduced aliphatic and aromatic species as well. [C/Fe] and [C/H] <1.0<-1.0 for a nearby Milky Way star is highly improbable (Hinkel et al., 2017; Lisse et al., 2017; Mészáros et al., 2025), and the aliphatic and PAH absorption features for 3I in the JWST NIRSPEC spectrum of Cordiner et al. (2025) look to be of reasonably normal strength.
Another hypothesis from the August 2025 observations suggests that 3I is in the same state of advanced thermal processing, mall loss, and age as analogously behaving 103P. Because of 103P’s very small size but moderate to high gas production rate, coupled with its lack of hypervolatile $\mathrm{CO}$ and presence of large $\mathrm{CO}_{2}$ -rich grains, it has been argued that it has been highly thermally processed by the Sun over hundreds to thousands of orbits, and is near the end of its life (Lisse et al., 2009). To get to this state, it has lost much of its original mass, so that its current surface used to be its core material, buried under kilometer’s worth of cometary material at its creation.
Another possibility is that we are observing a comet-like body from an exosystem with near-solar composition, but also a body with unusual alteration of its surface layers by Gyrs of continual galactic cosmic ray and stellar XUV irradiation kept at $T_{\mathrm{galactic\ ISR}}=15\mathrm{-}30\,\mathrm{K}$ (Zhu and Huang, 2014). This would produce an $\mathrm{CO}:\mathrm{CO}_{2}$ number ratio of $\approx 0.1$ if the surface material started out as $\mathrm{H_{2}O}:\mathrm{CO}_{2}$ ice only, without any starting $\mathrm{CO}$ (Maggiolo et al., 2025). However, if $\mathrm{CO}$ was initially present in the newly ejected comet’s surface at typical primordial-like abundance ratios of $\mathrm{CO}:\mathrm{CO}_{2}=0.5\mathrm{-}10$, the irradiated result should have a final $\mathrm{CO}:\mathrm{CO}_{2}$ ratio $\sim 1$ (Moore 1991). The observed SPHEREx $\mathrm{CO}:\mathrm{CO}_{2}$ ratio of $0.06\pm 0.03$ is consistent with a predicted value of $0.1$ for a starting pure $\mathrm{H_{2}O}:\mathrm{CO}_{2}$ ice case, and the large emitted dust grains are reminiscent of the mantle grains emitted from surface processed short-period solar system comets (Lisse, 2002) implying that 3I was thermally processed enough in its home system to severely deplete hypervolatiles in its near-surface layers before ejection and irradiation.
The way to distinguish between the [Complete thermal $\mathrm{CO}$ depletion of an object] versus [Surface $\mathrm{CO}$ thermal depletion followed by meters worth of galactic radiation treatment] hypotheses is to search for a marked change in the relative $\mathrm{CO}$ emission behavior of 3I  both in increased abundance and direct nuclear emission morphology, as it passes close to the Sun on 30 Oct 2025 and its interior regions, beneath its $\sim 10\,\mathrm{m}$ surface irradiation processed zone (Maggiolo et al., 2025), become heated.
V.2 ISO Evolution and Home System Implications
@@
A strongly thermally processed state for 3I has important implications. By analogy to 103P’s behavior, it tells us that 3I has undergone 100s to 1,000s of close passages by its natal host/birth star (we reject the possibility that 3I has undergone hundreds to thousands of close passages of random galactic stars on its passage to the solar system; this rate of close passage for a single object, even one as old as the galaxy, would imply a huge rate of interstellar object passages & detections in our solar system, which is not seen).
These close thermally altering passages could have been accomplished either as a primordial body formed in-place near its birth system’s $\mathrm{CO}_{2}$ ice line (to ensure minimal CO but extensive labile $\mathrm{CO}_{2}$ ice incorporation) and quickly ejected soon after PPD disk clearing, or, as seen currently in our own mature system, as an ice-rich KBO that became scattered into its inner system for a few hundred near-star orbits before ejection.
@@
Finally we end with a piece of scholarly speculation. Naive models of ISO creation would suggest the maximal number of ISOs created by a system should be when it hosts the most free-flying small planetesimals that are dynamically hot. I.e., when the system is very young, within the first few Myrs of the host stars formation. But this is also the era when a system’s protoplanetary disk is optically thick near its midplane and planetesimals in the densest part of the disk remain cold and thermally unprocessed. PPDs are seen to clear within $5\mathrm{-}10\,\mathrm{Myrs}$ of host star formation though (Rieke et al., 2005; Ben et al., 2025, and references therein), and the first heavy thermal processing of KBOs estimated to occur in $1\mathrm{-}40\,\mathrm{Myrs}$ (Davidsson, 2021; Steckloff et al., 2021; Lisse et al., 2021, 2022b; Parhi and Prialnik, 2023).
Planetary migration and Oort Cloud formation involving small body ejection processes can take 100’s of Myrs to complete (Brasser and Morbedeli, 2013; Garrod, 2019). We can thus plausibly argue that PPDs can clear and small body thermal evolutionary processes can occur more rapidly than dynamical processes can cool down and mature in a system, and we should see a mix of volatile rich, hypervolatile depleted, and totally devolatilized ejected objects as our ISOs. This is precisely what we have seen so far in highly-devolatilized 1I/Oumuamua, hypervolatile-rich 2I/Borisov, and well-thermally processed 3I.
@@ VI Conclusions and Future Work
In conclusion, the SPHEREx spacecraft observed ISO 3I from 01 to 15-Aug-2025 UT using 102 band, $R=40\mathrm{-}130$ spectrophotometry. Photometric imagery, spectroscopy, and light curves of the ISO were obtained. From these, robust detections of slightly extended water gas emission at $2.7\mathrm{-}2.8\,\mathrm{\mu m}$ and highly extended $^{12}\mathrm{CO}_{2}$ gas at $4.23\mathrm{-}4.27\,\mathrm{\mu m}$ were found. More tentative detections of $^{13}\mathrm{CO}_{2}$ and $\mathrm{CO}$ gas were seen. A slightly extended coma of $\mathrm{H_{2}O}$ was detected, and a huge surrounding atmosphere of $\mathrm{CO}_{2}$ extending out to at least 3’ was discovered. No jets were detected in the coma imagery, but a slight sunward pointing asymmetry was seen in the $\mathrm{CO}_{2}$ coma. Derived gas production rates for the 4 species were $Q_{\mathrm{H_{2}O}}=3.2\times 10^{26}\pm 20\,\mathrm{\%}$, $Q_{{}^{12}\mathrm{CO}_{2}}=1.6\times 10^{27}\pm 10\,\mathrm{\%}$, $Q_{{}^{13}\mathrm{CO}_{2}}=1.3\times 10^{25}\pm 25\,\mathrm{\%}$, and $Q_{\mathrm{CO}}=1.0\times 10^{26}\pm 25\,\mathrm{\%}$.
@@
Registration and co-adding of dozens of scattered light continuum images at $1.0\mathrm{-}1.5\,\mathrm{\mu m}$ produced a high S/N ratio image consistent with a stellar point source of no extension. The same images were used to create a lightcurve, for which variability $\lesssim 15\,\mathrm{\%}$ was seen from beginning to end of the observing period. The absolute brightness of 3I at $1.0\mathrm{-}1.5\,\mathrm{\mu m}$ is consistent with a nucleus of effective radius $R_{\mathrm{nuc}}\lesssim 2.5\,\mathrm{km}$, as reported by HST (Jewitt et al., 2025) surrounded by a 100 times brighter coma. The SPHEREx spectral continuum measured over 2 weeks is highly consistent with the continuum reported by JWST during its single visit to the object on 06 Aug 2025 (Cordiner et al., 2025). The $1.5\mathrm{-}4.0\,\mathrm{\mu m}$ continuum structure shows a strong feature commensurate with water ice + refractory organics absorption. The overall shape of this reflectance continuum is very similar to that reported for the cliff-type KBOs by (Pinilla-Alonso, 2025)
@@
The August SPHEREx observations provided the first detection of strong $\mathrm{CO}_{2}$ emission from the comet into a giant extended surrounding coma, proving that $\mathrm{CO}_{2}$ sublimation from large icy coma grains is the main driver for the observed outgassing activity and explaining its extended morphology as the product of $\mathrm{CO}_{2}$ gas sublimation that was undetectable in the first reported UVIS characterizations of the object. The moderate observed water gas production rate can be explained as an object just coming in from the cold depths of interstellar space that was observed at $r_{\mathrm{hel}}=3.2\,\mathrm{au}$ from the Sun outside the water ice line.
@@
In many ways the observed activity behavior of 3I, including its preponderance of $\mathrm{CO}_{2}$ emission, lack of $\mathrm{CO}$ output, small size, and predominance of large chunks of water-ice rich icy material in a flux-dominant coma is reminiscent of the behavior of short period comet 103P, target of the NASA Deep Impact extended mission in 2010 and a “hyperactive comet” (A’Hearn et al., 2011; Lisse et al., 2009). Such an object could have undergone hundreds to thousands of close perihelion passages by its natal birth star inside the system’s water ice line, as 103P has by our Sun. Whether this was due to 3I forming within a few au of its birth star or being scattered into this region after formation but before ejection is not clear.
@@
Compared to the naïve expectation that solar systems shed the majority of their comets in their first few Myrs of existence when their small icy planetesimals are most abundant, most hypervolatile ice rich, and most dynamically hot, so that all ISOs should be like 2I/Borisov, this suggests that ISOs are thermally processed quicker than they dynamically relax. This hypothesis needs to be tested by population studies of many more ISOs than the 3 currently known, and hopefully many more will be studied by the SPHEREx mission’s all-sky survey.
@@
This paper relies on data obtained by the SPHEREx Observatory (operated by Caltech and JPL on behalf of the National Aeronautics Space Administration under contract 80GSFC18C0011) that is hosted by IPAC as part of the IRSA archive (10.26131/IRSA652 SPHEREx Quick Release Spectral Images - QR2). We are indebted to the SPHEREx project for obtaining additional pointings at 3I, approximately 100 percent more, than their nominal survey would have obtained less than 3 months after the start of the survey by testing and implementing a new commanding protocol. We have also used optical lightcurve trending data kindly supplied by S. Yoshida. The research was carried out at the Jet Propulsion Laboratory, California Institute of Technology, the Johns Hopkins Applied Physics Laboratory, and Arizona State University under a contract with the National Aeronautics and Space Administration (80NM0018D0004). The authors also acknowledge the Texas Advanced Computing Center (TACC)  [^10] http://www.tacc.utexas.edu at The University of Texas at Austin for providing computational resources that have contributed to the research results reported within this paper.